OpenStudy (xobriley):

3.Which property is illustrated by the statement? 3+5x = 5x+3 *a) commutative property of addition b) associative property c) distributive property

OpenStudy (xobriley):

@gabbyalicorn

OpenStudy (eboogie4):

I think distributive cause I did that b4 but I would get a second opinion to b sure

OpenStudy (xobriley):

Thanks

OpenStudy (eboogie4):

yep np if u ever need help I will always try :)

OpenStudy (gabbyalicorn):

This might help for future questions like this. 3(x + 2) – 4x original (given) statement 3x + 3×2 – 4x Distributive Property 3x + 6 – 4x simplification (3×2 = 6) 3x – 4x + 6 Commutative Property (3x – 4x) + 6 Associative Property

OpenStudy (xobriley):

I was wrong

OpenStudy (gabbyalicorn):

Hmm?

OpenStudy (eboogie4):

what do mean?

563blackghost (563blackghost):

I would say you are correct... \(\Huge{a+b=b+a}\) That is the format of your question which is fact communicative property...

OpenStudy (xobriley):

WAS I WRONG

563blackghost (563blackghost):

No you are not wrong @XOBriley